OpenStudy (anonymous):

In a recent year the total sales at a restraunt A were $325,000 more than restraunt B. Total sales at the two stores were $625,000,000what were the sales at for the yearat each store?

OpenStudy (anonymous):

A= 325000+ B A+ B= 625000

OpenStudy (anonymous):

Solve for A and B

OpenStudy (anonymous):

but how do you do it

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

A-B= 325000 A+ B= 625000

OpenStudy (anonymous):

Add the two equations

OpenStudy (anonymous):

wait add 325000+625000000 it has 6 zeros

OpenStudy (anonymous):

2 A = 950,000 A= 475000

OpenStudy (anonymous):

so which one is A and B

OpenStudy (anonymous):

A+ B= 625000 B = 625000-A= 625000-475000=150000

OpenStudy (anonymous):

A restaurant A B restaurant B

OpenStudy (anonymous):

A= 475,000 B= 150,000

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

YW